OpenStudy (anonymous):

what is the simplified form of 1/5xy * 10x^2/7y

OpenStudy (anonymous):

\[\large \frac {1}{5xy}\times \frac {10x^2}{7y}\]

OpenStudy (anonymous):

that your problem?

OpenStudy (anonymous):

yep